OpenStudy (anonymous):

\[((1/y^2)-(3/y^4)) * (y +5y^3)\] differentiate

OpenStudy (thomas5267):

Some \(\LaTeX\) tips: \times=\(\times\) \frac{a}{b}=\(\frac{a}{b}\) y^{-2}=\(y^{-2}\). The braces are used to group things together. Okay, so do you have to differentiate this with respect to y? \[ (y^{-2}-3y^{-4})(y+5y^3) \] I think negative indices are easier to work with when doing calculus.

OpenStudy (anonymous):

yes i do have to differentiate with respect to y. i understand what you but now do i have to do the product rule?

OpenStudy (thomas5267):

I wouldn't do product rule. I would expand those bracket and differentiate term by term.

OpenStudy (anonymous):

ohh ok i understand now thank you very much that makes everything much simpler

OpenStudy (thomas5267):

You are welcome.